These questions require application of economic theory relating to elasticity of demand and supply. All calculations must be shown in full. Answer ALL the questions.

Q.3.1 A store that sells rice discovers that when the price of 1kg rice Is R24 per kilogram, the quantity demanded is 306 kgs per week. When the price decreases to R21 per kg, then the sales increase to 340 kgs per week. Use this information to answer questions

Q.3.1.1 and Q.3.1.2 below. Q.3.1.1 Determine the price elasticity of rice using the Arc method.

Q.3.1.2 Discuss the relationship between the price elasticity of rice and the total revenue the store received from the sales. Advise the store on an appropriate pricing strategy.



1
Expert's answer
2021-05-26T15:33:23-0400

1.1

"Arc_{(Ed)}=\frac{\frac{Q_2-Q_1}{(Q_2+Q_1)\div2}}{\frac{P_2-P_1}{(P_2+P_1)\div2}}"


"=\frac{\frac{340-306}{(340+306)\div2}}{\frac{21-24}{(21+24)\div2}}"


"=\frac{34}{23}\div\frac{3}{22.5}"


"=-0.789"


1.2

The price elasticity of demand is inelastic meaning an increase in price will lead to a decrease in total revenue. The a appropriate pricing stategy willl be competitive pricing strategy whereby price setting is based on what the competition charges. For example charging low price will increase the total revenue.